### v1

# Not Everyone Wins with LLMs: Behavioral Patterns and Pedagogical Implications in AI-assisted Data Analysis

Abstract.
LLMs promise to democratize technical work in complex domains like programmatic data analysis, but not everyone benefits equally. We study how students with varied expertise use LLMs to complete Python-based data analysis in computational notebooks in a non-major course. Drawing on homework logs, recordings, and surveys from 36 students, we ask: Which expertise matters most, and how does it shape AI use? Our mixed-methods analysis shows that technical expertise – not AI familiarity or communication skills – remains a significant predictor of success. Students also vary widely in how they leverage LLMs, struggling at stages of forming intent, expressing inputs, interpreting outputs, and assessing results. We identify success and failure behaviors, such as providing context or decomposing prompts, that distinguish effective use. These findings inform AI literacy interventions, highlighting that lightweight demonstrations improve surface fluency but are insufficient; deeper training and scaffolds are needed to cultivate resilient AI use skills.
large language model, data science assistant, data analysis, AI coding assistant, human-AI interaction
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXbooktitle: 2025 Preprintprice: 15.00isbn: 978-X-XXXX-XXXX-X/YY/MMccs: Human-centered computing Interactive systems and toolsccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generation
1. Introduction
Large language models (LLMs) hold the promise of democratizing complex technical tasks such as data analysis by assisting users in programming, analysis, and modeling (Ma et al., 2025; Pickering et al., 2025). Despite the grand vision, we continue to see mixed results in whether and how people benefit from AI (Paradis et al., 2025). There have been conflicting reports on whether AI agents help developers improve programming efficiency (Becker et al., 2025; Song et al., 2024); only some novices seem to perceive reduced barriers (Pickering et al., 2025; Kazemitabaar et al., 2023). In fact, performance often hinges on users’ prior expertise, ranging from domain knowledge to AI literacy (Chen et al., 2024; Nguyen et al., 2024a; Zi et al., 2025). For instance, some work argues that less experienced developers benefit more from AI assistants (Liang et al., 2024; Peng et al., 2023), and those with basic AI literacy may avoid common pitfalls in prompting or interpreting model outputs (Puppart and Aru, 2025; Annapureddy et al., 2025). However, these findings remain fragmented across different populations and tasks, making it challenging to weigh which types of expertise matter most (e.g., whether we should focus on teaching domain knowledge or AI usage knowledge). Moreover, most prior analyses measure only outcomes (e.g., task success) and are conducted in controlled, short-term lab settings (Kazemitabaar et al., 2024a; Pickering et al., 2025; Zheng, 2023). As a result, we have yet to demystify how users’ AI interaction behaviors get shaped by their (lack of) certain expertise in real-world workflows.
In this work, we aim to examine how varied forms of expertise influence AI use and outcomes in a naturalistic classroom setting: a graduate-level course, Data Science for Product Management (DSPM). The class offered a unique opportunity to study LLM-supported workflows: it included a diverse cohort of 40 students from CS majors to business students transitioning into technical roles, allowing us to examine varying domain expertise, LLM literacy, and communication skills. Students worked independently on a series of data analysis assignments (e.g., Data Cleaning) in Google Colab, had full access to the built-in Gemini LLM assistant (Figure 1), and their activity was captured through logs, surveys, and screen recordings. Through this setup, we investigate three questions of increasing depth:
RQ1 (Section 5): Do LLMs close the performance gap between students with different expertise? We first conduct quantitative analysis on the relationship between students’ self-perceived expertise (in technical, LLM, and communication domains) and their grades across four assignments. Despite all students having equal access to Gemini, we find that self-reported technical expertise significantly predicted student grades, while LLM expertise and communication skills did not. This suggests that LLMs do not fully close the expertise gap — students with stronger programming and data science backgrounds outperform their peers when using AI. Many dimensions of expertise were highly correlated in our study, making it difficult to isolate their independent contributions. This indicates that fully isolating expertise may require more deliberate sampling and experimental control in future work.
RQ2 (Section 6): How do differences in expertise translate to AI usage behaviors? We analyze students’ interaction logs using a novel behavior-grounded codebook, developed via thematic analysis and scaled up through LLM-based annotation. Our method segments over 7,000 raw log events into intent-driven episodes and steps, where each step is annotated with the presence (or absence) of effective AI behaviors, such as, decomposing tasks, providing contexts, and using AI for explanation or critique. Our analysis shows that technical experts use AI more strategically. They write clearer prompts, are more likely to provide necessary context, and use AI proactively to explore alternative solutions or improve plans (e.g., asking for better visualizations). In contrast, novices rely more on AI to resolve immediate challenges, such as debugging errors or explaining confusing outputs. These behavioral patterns offer a more nuanced explanation of why expertise continues to matter, even in AI-supported environments.
RQ3 (Section 7): What and how can we train students on AI usage? We analyze missed opportunities for AI use by identifying log steps where AI could have been helpful but was not used. From this, we distilled a set of AI-use competencies spanning across four knowledge dimensions Conceptual, Procedural, Metacognitive, Dispositional. We show that some AI-use skills can be improved via live coding demonstration and extended time-on-task (e.g., regarding prompt quality), while others likely require guided practice and feedback (e.g., regarding evaluation), and are hard to cover via demonstration (e.g., metacognitive and dispositional skills). In sum, we make three main contributions:
  • Empirical evidence on expertise vs. AI use: We show that even in open-ended, LLM-supported workflows, technical expertise remains a significant predictor of success, and self-reported communication skill and AI familiarity alone do not suffice.
  • Complex log-based characterization of AI use: We introduce a behavior-oriented, LLM-based annotation schema that segments over 7,000 raw log events into semantically meaningful user stories. Compared to prior work that analyzes linear chat logs (Lee et al., 2022), our setup in Google Colab involves multiple LLM-powered functions with different affordances and model access, resulting in more nuanced observations about when, how, and why students turn to AI support (e.g., comparing behaviors between AI-as-chat versus AI-as-inline-tool). We open-source all of our scripts, including log data collection, cleaning, annotation, and analysis: https://github.com/mqo00/dspm
  • Pedagogical implications: We identify which AI skills can be improved via lightweight instruction vs. require deeper pedagogical investment. This informs future curricula and tool design for training effective AI collaborators.
2. Related Works
2.1. Novice versus Expert Strategies in Programming with AI
HCI and educational research highlight systematic differences between novices and experts when collaborating with AI for programming and modeling. In agent-based modeling with ChatGPT, experts benefit more than novices and rely on stronger skills in task decomposition and evaluation of AI outputs (Chen et al., 2024). In data-science code comprehension, experts inspect the data deeply, verify assumptions about transformations, and navigate outputs with clear goals—behaviors novices underuse (Lum et al., 2025). At the same time, studies document failure patterns such as “prompting rabbit holes,” where users iterate on prompts without convergence (Tie et al., 2024). However, most prior studies also focus on challenges brought by domain expertise differences, such as experts verifying and examining the data more deeply, and few focus on domain-agnostic LLM use patterns or skills regardless of expertise (Section 2.3). Together, these findings motivate our research questions: measure whether LLMs truly reduce expertise gaps, and analyze which behaviors distinguish effective from ineffective AI use in authentic notebook workflows.
2.2. Interfaces and Log Mining for AI-Assisted Data Analysis
AI-based computational notebook assistants have investigated a vast design space of scaffolding users’ interactions with LLMs during data analysis (Mcnutt et al., 2023). For example, prior works have explored providing phasewise or stepwise decomposition interfaces (Kazemitabaar et al., 2024a), structuring how users express intent and context with ephemeral UI (Cheng et al., 2024), manipulating intent (Kim, 2025), and embedding explanations and documentation support (Wang et al., 2022). These explorations brought different values to streamline the AI-assisted data analysis tasks, such as improving users’ perceived control (Kazemitabaar et al., 2024a) and lowering interaction barriers for novices (Wang et al., 2022; Cheng et al., 2024). However, they do not necessarily improve objective performance measures such as task success or completion time, as UI scaffolds on their own do not ensure learning outcomes or allow us to analyze the AI-use skills students need.
Beyond interface design, many efforts have explored mining logs of user interactions to derive useful insights. For example, recent works have explored logging and learning analytics in Jupyter notebooks to track students’ interactions with a chatbot AI (Valle Torre et al., 2025) or log novices’ data science workflow (Zhao et al., 2024). These approaches highlight the promise of leveraging naturally occurring clickstreams to study novice reasoning at scale (Claypool et al., 2001; Hijikata, 2004; Huang et al., 2012; Chen et al., 2025b), but also the challenges: raw logs are noisy, fragmented, and difficult to interpret without additional semantic structure e.g., a long pause may indicate either confusion or deep focus. To enable higher-level reasoning, we take inspiration from recent work, to extract more semantically meaningful insights from the raw logs by utilizing LLMs’ reasoning and Theory-of-Mind capabilities for reasoning process (Viswanathan et al., 2024; Rabinowitz et al., 2018; Kosinski, 2024; Strachan et al., 2024). To this end, we propose to combine log-based methods with behavior-grounded coding to identify success and failure patterns in AI-assisted notebook tasks, and identifying needed AI skills from the analysis. By semantically enriching logs into user story episodes and mapping them to concrete challenges, we aim to advance understanding of why some learners succeed while others fail, and what kinds of competencies need to be taught (Section 7.1, 7.2).
Figure 1. Google Colab notebook environment with embedded Gemini assistant. Students could interact with Gemini in two main ways: code cell generations (generateCode) and a conversational side chatbot (converse). There are also additional buttons that support error explanation, code explanation, and visualization generation.
2.3. Training for AI-Assisted Data Analysis and Effective LLM Use
Prior works have also explored instructions specifically for AI-assisted data analysis tasks. For example, Zheng (2023) allowed ChatGPT use in a graduate data science task and found students reported that AI was more helpful for explaining code and less for open-ended problem-solving that needs context, especially on LLM-hard tasks where AI has limited capacities. LLM-enhanced feedback for an R-based data science tutoring system was also found to improve students’ performance, while the verbosity of AI explanations sometimes caused cognitive overload (Letteri and Vittorini, 2025). Nonetheless, these works mostly focus on analyzing surveys and prompt exchanges, without behavioral log data that is important for capturing the full experiences and interactions with LLM (Kazemitabaar et al., 2024b). Moreover, these studies may not represent the most authentic workflows of how students would use LLM for computational data analysis, when there are popular tools such as Google Colab (Google, 2025) where AI is embedded in the notebook environment.
There have also been some works that focus on task-agnostic, domain-general AI-use skills, such as prompt engineering. However, students do not learn from code feedback and self-experimentation (Nguyen et al., 2024b); teaching specific prompt tips also may not be very effective at influencing students’ real prompting approaches and success, and also risks getting obsolete with model updates (Ma et al., 2025). Some prior works have shown promise in teaching requirement specification (Ma et al., 2025) or requirement elicitation (Lojo et al., 2025) for LLMs. However, no work has comprehensively mapped AI-use competencies across the full workflow or disentangled which skills are teachable via simple demonstrations versus guided practice and targeted feedback. Our study fills this gap by linking fine-grained behaviors to outcomes and by quantifying post-demo shifts in skill acquisition.
3. Course Structure and Task Setup
3.1. Course Context
Our IRB-approved study was conducted in Spring 2025 in Data Science for Product Managers (DSPM) at an R1 University in the United States. The course serves students with highly diverse academic and professional backgrounds, including product managers, engineers, and graduates of non-technical disciplines. This diversity creates a pedagogical tension: assignments that are trivial for some are overwhelming for others. The setting therefore provides an ideal opportunity to examine whether and how LLM support can reduce expertise gaps in authentic workflows.
The class included four sequential assignments (HW1-HW4), each covering a primary data analysis stage: (1) data cleaning, (2) exploratory data analysis (EDA), (3) lightweight machine learning, and (4) data-driven storytelling. All assignments used the same dataset (more details in Section 3.2), enabling continuity across tasks. Lectures covered corresponding data science topics and basic GenAI use with a demo (more details in Section 7.1).
Students worked in the Google Colab notebook with access to the built-in Gemini assistant (Google, 2025) [^1] We chose Google Colab instead of Jupyter Notebook because it is easier to set up for students who do not have any programming background, and Colab has an in-house LLM (Gemini) that is free to use and provides various affordances. Specific model used by Google Colab in Spring 2025 is unknown, but it could be gemini-2.0-flash or gemini-1.5-flash. We discuss the limitations of this approach compared to building on the open-source Jupyter Notebook in Section 8.2., but external GenAI systems such as ChatGPT were disallowed. Colab Gemini offered five interaction modes, including a chatbot, code cell suggestions, and buttons for error explanations, code explanations, and visualization generation with a dataframe (Figure 1 shows two of the main interfaces used for code generation: chatbot and code cell).
Table 1. An overview and comparison of the homework setup, dataset, and rubrics. Overall task: analyze the data to gather some insights to present and guide product development across four stages of analysis (data cleaning, EDA, ML, storytelling).:
HW0 Task HW1-4 Task
Dataset Car dataset with 649 entries (crowdsourced from the prior semester of this class). Columns: Car Make, Model, Dealer/Individual, Price, Year, Location, Mileage, Doors, VIN# (9 in total). Game dataset with 1360 entries (adapted from a Kaggle dataset, with renamed fields and inserted dirty data). Columns: name, description, developer, rating, genre, released_on, #total_players, #current_players, #saves, #listed, #reviews, selected_reviews (12 in total).
Cover Story You’re a product manager of Honda Accord, and the company is aiming to launch an online car price valuation tool to help buyers and sellers of used and new Honda Accord cars. Your team has collected a dataset for you by manually searching for 2008 to 2022 Honda Accords for sale in the U.S. You’re a product manager of Steam, a video game distribution platform, and the company is aiming to launch an analytics platform for game developers that provides insights into market trends. Your team has collected a dataset for you by manually scraping some popular games on Steam since 2000.
Template,
Timing &
Tasks
[Uncaptioned image] [Uncaptioned image]
(screenshot of HW0 template for T1 data cleaning) (screenshot of HW2 template for T2 EDA)
In-lecture activity. Four tasks (T1–T4), each $\sim$15 minutes. Half of the students use AI and half do not, counterbalanced across two tasks. Bi-weekly assignments (HW1–HW4), each structured around T1–T4, with extended time for completion and free use of AI in Google Colab.
Rubrics T1 Data Cleaning (max 15 pts): Clean at least five columns of the dataset (e.g., duplicates, missing values, outliers, types, formats). Rubric (for each of the 5 columns) — 0: no cleaning, 1: very incomplete, 2: partial, 3: complete.
T2 EDA (max 45 pts): Ask and answer meaningful data questions using stats/visualizations. Rubric (for each of the 5 questions/answers/processes) — 0: missing, 1: simple/vague, 2: partial, 3: advanced/complete.
T3 ML (max 15 pts): Implement and document an ML workflow (train/test split, build model, measure performance, compare models, decide best model). Rubric (for each of the 5 steps) — 0: not implemented, 1: very incomplete, 2: partial, 3: complete/correct.
T4 Storytelling (max 15 pts): Produce a clear data-driven report with framing, stats/modeling, visualization, storytelling, and data discussion. Rubric (for each of the 5 aspects) — 0: missing, 1: very incomplete, 2: partial, 3: complete/clear.
3.2. Homework Design and Procedure
Assignments were distributed as Google Colab notebook templates with minimal instructions and starter code, requiring students to “fill in the blank” using Python.
Students first completed a diagnostic HW0 during lecture to familiarize themselves with the tools and task sequence. HW0 covered all four tasks but used a different dataset than HW1-HW4 (details in Table 1). We asked students to spend about 15 minutes on each task, where half the class completed the first two or last two tasks without using GenAI, providing a baseline for comparison. Students then completed HW1–HW4 over the semester on a bi-weekly basis, each focused on a different aforementioned task. This setup enabled us to probe both short-term performance and longer-term workflow strategies, and to examine whether the LLM helps close the expertise gap. Details regarding the dataset, task, and rubrics are presented in Table 1.
For each assignment, students submitted (1) a completed notebook, (2) exported log files of interactions, and (3) a post-survey reflecting on challenges and AI usage (details in Section 4.2 and Appendix B). For HW0, students recorded their screen via Zoom, creating paired video–log traces for deep qualitative analysis. For HW1–HW4, bonus credit was offered for optional 60-minute think-alouds facilitated by one author, during which students would verbalize their thoughts while doing homework as usual.
The assignments were graded by two teaching assistants (TAs) for the class. The TAs independently graded 20% of the student submissions in HW0 and achieved a Cohen’s kappa $\kappa=0.71$, which suggests substantial inter-rater reliability (McHugh, 2012). Weekly calibration meetings were held to align rubric interpretation.
4. Method
4.1. Participants
Figure 2. Distribution of self-rated technical expertise, LLM expertise, and communication expertise scores.
A total of 41 students completed HW0, with a slight attrition to 36 students by HW4. Students came from a wide range of fields, including finance, telecommunications, journalism, physics, music, and computer science. This diversity highlights both the promise and the challenge of designing AI supports that work for “everyone.”
We collected pre-survey data on different self-reported expertise, including implementation (e.g., Python proficiency), domain (e.g., data science expertise), toolkit (e.g., Colab/Jupyter Notebook, pandas, and altair, which are the specific environment and libraries used for our tasks), model (e.g., GenAI/LLM proficiency), and non-technical skills (e.g, communication abilities). We initially aimed to analyze whether these factors influence effective LLM use in different ways; however, in our class, some skill areas were moderately correlated (e.g., a Spearman correlation $\rho=0.52$ for Python and data science items, and $\rho=0.72$ between data science and tookit items), making it difficult to disentangle their independent contributions (we discuss more about future work in Section 8.2).
Therefore, we operationalize students’ technical expertise by grouping the survey answers to programming, data science, and toolkit expertise questions, such as “How many years of experience do you have in using the following tools? [Python]” and “How confident are you in your ability to do the following tasks on your own? [Exploratory data analysis]” (more details in Appendix C). Similarly, we merged highly correlated items into composite measures of LLM expertise (e.g., “How would you assess your level of expertise in the following areas? [GenAI/LLM]”) and communication expertise (e.g., “I am confident in my communication ability in general.”). Items that were highly skewed toward zero or not informative were excluded. The distribution of these expertise among students is shown in Figure 2, where students were classified as novices or experts using a median split of self-ratings.
4.2. Data Collection
Our dataset combines homework submissions, interaction logs, surveys, and screen recordings; for each student, we collected:
  1. (1)
    Notebooks and grades (HW0-HW4 submissions; rubric-based),
  2. (2)
    Behavioral logs (HW0-HW4; code edits, executions, errors, prompts, model responses, timestamps, etc.),
  3. (3)
    Surveys (one pre-course survey, five post-HW surveys, one post-course survey)
  4. (4)
    Screen recordings (HW0 for all students; HW1–HW4 for think-aloud volunteers)
Surveys
We administered structured surveys at multiple points in the course (see Appendix B for full questions):
  • Pre-Survey (start of semester): Collected demographics, academic and professional backgrounds, Likert-scale questions on self-perceived expertise, interest, confidence in personal vs. GenAI ability in different fields, etc.
  • HW Post-Surveys (after HW0-HW4): Captured task-specific experiences. Each survey asked which GenAI tools were used (e.g., Colab Gemini, ChatGPT), perceived task difficulty, effort, frustration, engagement, self-assessed success and confidence, and amount of peer help. Open-ended questions probed challenges with GenAI (and without GenAI in HW0), how GenAI was used, its helpfulness, and what could improve performance.
  • Final Post-Survey (end of semester): Collected self-assessed expertise, interest, and confidence across domains (similar to pre-survey). Probed on similarities and differences across tasks, what they learned, and what they still wanted to learn.
   "event_id": "evt_67",
   "hw": "hw0",
   "studentid": [student id],
   "timestamp": 1741801351065,
   "time_since_started": 3642517,
   "event_category": "prompt",
   "event_type": "converse",
   "event_name": "enter_prompt",
   "event_content": "What are the prices for Honda
                     cars manufactured after the year 2015",
   "task": "t2_eda",
   "converse_chatid": "chat5",
   "turn": "3",
   "event_response": [LLM Output]
Figure 3. An example log event format.
Behavioral Logs
We deployed a custom instrumentation pipeline in Colab to capture both notebook interactions (code executions/errors), selected UI events (pressing explain-code/error buttons) and LLM use (prompts/responses). Students installed a Tampermonkey script that intercepted requests to Colab’s converse and generateCode endpoints and captured clickstream events. The script added a “Download History” button to the Colab toolbar and stored code cells, chat messages, event metadata, and responses to browser localStorage, which students exported as JSON files upon submission (or auto-save when the storage is full).
In total, we collected 16,315 logged events (see example in Figure 3) and 44 hours of screen recordings across five assignments. Logs captured a rich sequence of AI interactions, code edits, and executions, supporting both quantitative distributional analyses and qualitative episode-based coding.
4.3. Data Analysis
Throughout the paper, we conducted a mixture of quantitative and qualitative analyses to answer different research questions. Here, we only provide an overview of the analysis methods, and we detail the method used to answer each research question in its corresponding sections:
  • RQ1: To assess whether different expertise makes a significant difference in grades (Section 5), we conduct regression testing on student homework data and competency scores;
  • RQ2: To characterize students’ AI usage behavior in relation to their task completion sequences, competency, and grades (Section 6), we perform thematic analysis identifying key AI behavior types, use them to annotate student logs, and then quantitatively analyze the resulting distributions;
  • RQ3: To further identify skills essential for effective AI usage (Section 7), we qualitatively analyzed survey responses as well as the aforementioned student log behaviors, identifying skill challenges and learning opportunities across conceptual, procedural, dispositional, and metacognitive dimensions.
5. RQ1: Do LLMs close different expertise gaps?
Analysis method: Quantitative regression modeling.
We assess whether different expertise (Figure 2, details about expertise score calculation in Section 4.1) makes a significant difference in grades in HW1–HW4. We fit a linear mixed-effects regression model with grades (normalized scores) as the outcome and technical, LLM, and communication expertise scores as predictors, including random intercepts for participant and assignment (Bates et al., 2015).
Table 2. Linear mixed-effects regression predicting homework performance from different expertise. Participant and assignment were used as random intercepts.
Predictor Estimate SE df p
Intercept 89.88 17.84 33.82 $001^{***}$
Technical 6.09 2.86 30.40 $041^{*}$
LLM 3.05-3.05 2.69 30.77 .266
Commun. 2.54-2.54 2.85 30.97 .381
Analysis results: Technical experts can use LLMs more effectively than others.
As shown in (Table 2), only technical expertise significantly predicted higher grades ($\beta=6.09,p=.041$). LLM and communication expertise had no significant effects and did not predict grades. This suggests that even with equal access to the Colab Gemini assistant, students with stronger programming and data science backgrounds continued to outperform peers.
Note that students in our class had a median of 4/5 for their communication expertise (see Figure 2), which means that the communication factor may not have a lot of variance to predict grades. Meanwhile, students’ median of LLM expertise is 3.67/5, but everyday use of LLMs or their self-rated LLM knowledge may not straightforwardly transfer to complex programming data science tasks. Additionally, the expertise scores here are all operationalized as self-reported survey items, and there are some known issues in miscalibration of self-perceived expertise (Kruger and Dunning, 1999). Future work should explore more rigorous ways to measure expertise and recruit a more diverse set of students (more in Section 8.2).
Logs and surveys showed that students engaged substantially with AI during coursework: 15% of their logged events were prompt-related, and homework post-surveys all confirmed that students spent non-trivial amounts of time with AI tools. Despite substantial AI usage, the technical expertise gap remained — technical experts seem to benefit more from LLM use than novices, which is consistent with some prior literature (Chen et al., 2024).
6. RQ2: How do differences in expertise translate to AI usage behaviors?
The analysis in RQ1 showed that students’ self-perceived technical expertise continued to predict performance even when all students had access to the Colab Gemini assistant. Here, we dive into student log data to examine why: how students use AI during workflows, and how it relates to different levels of technical expertise?
6.1. Analysis method: Thematic analysis and LLM-automated annotation
6.1.1. Data filtering.
As mentioned in Section 3, we required students to use the built-in Gemini and avoid using other LLMs. We added the precaution of asking students to self-report their additional LLM model use, if any. Despite our best efforts, we still noticed students using external AI tools, which tend to result in reasonable grades but no or few user logs captured, or logs with abnormal behaviors, e.g., copy-pasted large blocks of code with unfounded originality could not be found.
To ensure data quality for log analysis, we excluded these cases and retained logs from 28 students (16 experts [^2] The expert vs. novice for the rest of the paper are with respect to technical expertise., 12 novices) across 36 submissions (20 from experts, 16 from novices) across HW0, 1, and 2 (7,315 logged events). Specifically, we filtered out students who self-reported other tool use like ChatGPT or displayed other tool use in recording (if available), as well as those who disengaged or spent too much time working on tasks without AI, reflected as log length being less than the median. We also excluded logs from HW3-4 to enable meaningful comparison with HW0, where most students did not attempt Task 3 and 4.
Figure 4. (A) Average frequency of AI use codes per student (refer to Table 3 for code definitions), aggregated across all steps and clustered by AI usage code types, including improve, code, evaluate, explain, and prompt (quality). (B) Example annotated log snippet. The student used AI in two actions (pressing a get visualizations button in Colab and entering a chat prompt “How do I clean my dataset”), reflecting AI usage ai_suggest_steps_or_plan and ai_explain_concepts and indicating that the student formed the intent of the first data cleaning episode.
6.1.2. Semantically meaningful log analysis.
Raw logs consist of fragmented sequences of actions that lack meaningful semantics. To recover actual student usage and rationale, we analyze user logs through a hierarchy of episodes, steps, and log actions (as shown in Figure 4B).
Specifically, we segmented logs into episodes — goal-driven units of activity (e.g., ”remove duplicates in data column X”). Inspired by Norman’s Stages of Action for user cognitive processes (Tenner, 2015), we further broke each episode into four steps — forming intent, expressing the intent as an input (e.g., writing code or prompt), obtaining and understanding output (e.g. inspecting the code execution results or errors), and assessing output (e.g., deciding if the intent is sufficiently fulfilled). Each step is achieved through a sequence of captured log actions. Within each step, we annotated whether the student encountered challenges, whether the step succeeded or failed, what AI usage behavior occurred, and what additional AI usage behavior could have helped, or missed opportunities of benefiting from more effective AI usage.
Develop the codebook.
To answer “how are users using LLMs”, we identify the core AI usage behaviors by developing a codebook (Table 3). Two authors thematically analyzed a subset of homework logs and screen recordings, following a hybrid of deductive and inductive thematic analysis approach (Braun and Clarke, 2006; Fereday and Muir-Cochrane, 2006). To maximize verifiability, we used five logs and screen recordings from HW0 and five logs and think-aloud recordings from HW1-4. The AI code types in our codebook included both the AI use objectives (improving, coding, evaluating, explaining), and quality of the corresponding student prompts, as using an AI is not the same as using the AI well (echoing observations in literature (Zamfirescu-Pereira et al., 2023; Ma et al., 2025)).
Table 3. Codebook for AI usage behaviors. Code reflect different usage objectives (to have AI improve plans, assist coding, explain concepts, or evaluate deliverables), as well as usage quality (whether prompt was clear enough, etc.). For the prompt quality category, we provide negative examples since other rows already demonstrate positive examples.
Code Definition and Examples
Improve (Plan)
ai_breakdown_intent Decompose a complex goal into manageable pieces to form the correct task intent. E.g., “How to improve performance of model. give step by step.”
ai_improve_prompt Refine the wording of a prompt for clarity, specificity, and completeness before generation. E.g., “plz improve my prompt: [prompt]”
ai_suggest_steps_or_plan Provide a step-by-step plan or workflow (e.g., data cleaning $\rightarrow$ modeling $\rightarrow$ evaluation) for the task. E.g., “How to do data cleaning?” or ”How to build ML models?”
Code
ai_generate_code Produce code that implements a clearly stated action to achieve the intended output. e.g., “Write the code to calculate the distribution of the cars by location.”
ai_edit_partial_code Modify only a selected snippet or function instead of rewriting the whole cell or file. e.g., [Ask AI to only edit part of the code instead of regenerating the entire cell.]
Explain
ai_explain_bug_or_error Explain an error or traceback and outline a concrete fix. E.g., Copy-paste error message and use ”explain error” to understand and fix it.
ai_explain_code_or_api Interpret code or clarify specific APIs (e.g., describe what a function does and how to use it). E.g., “Explain what pd.get_dummies does”
ai_explain_concepts Provide plain-language explanations of data science or machine learning concepts. E.g., “How do data cleaning steps work?”
Eval
ai_critique_output Evaluate results or insights for correctness, alignment with goals, and suggest improvements. E.g., “Check my insight and provide improvements”, “Is it a good model to work with?”
Prompt (Quality)
clear_instruction_in_prompt Write prompts with explicit action verbs and desired outcomes (avoiding vague instructions like “fix it”). “Fix it” without saying what ”it” is would be unclear.
decompose_task_in_prompt State the problem as ordered sub-tasks or “step-by-step” instructions within the prompt. E.g., “This was my result array [copied output]” without intent would be too vague.
provide_context_in_prompt Include necessary details (e.g., dataframe names, column names, prior steps) and avoid vague pronouns to prevent misinterpretation. E.g., “Don’t make up non-existent column name” without providing actual context does not help resolve hallucination.
Scale up with automated log annotation
Because manual annotation was time-intensive [^3] Even with the video and with one author already familiar with the session, on average, it took ¿ two hours to go through each log. It is even more difficult to decipher logs that do not come with video recordings, due to, e.g., readability issues (certain log events would contain more notebook context information with code cell metadata that are hard to manually read through)., we scaled up by using an LLM to label log steps with the codebook (claude-sonnet-4-20250514 for its reasoning capability, with temperature=0 for deterministic generations, full prompts in Appendix D). We had the model identify the challenges as well as AI usage — either the student used AI (usedused), or the student missed an AI use opportunity (missedmissed) [^4] Our complete annotation output also contains a series of other features we used for supporting the model’s inference time reasoning procedure and the later-on human verification, e.g., log_id, the reasoning behind certain label, etc. Complete prompt with the output format and other hyperparameters are in Appendix D..
Validation of LLM annotation.
Our annotation produced, on average, 7.6 episodes per log file, with 5.6 steps per episode. Out of a total of 7315 log events, there are 1483 steps (form intent, write code, understand output, and assess output) across all episodes, and 920 (62.0%) of them are annotated with AI usage ($used+missed$). The distribution of average AI usage frequencies by all the codes is shown in Figure 4. An author verified 10% of the log annotations, and LLM achieved an accuracy of 75% (specifically, 81.3% accuracy for step success, 76.7% for encountered challenge, and 70.9% for AI use behaviors annotations).
6.2. Analysis results: AI usage change across steps and expertise
Figure 5. Distribution of the AI usage (both usedused and missedmissed) aggregated by code type and episode steps.
Steps vs. AI usage: AI can help throughout stages, with clear mapping of “how”.
As shown in Figure 5, there is a clear relation between how AI is used and when. Behaviors associated with AI improving (e.g., ai_suggest_steps_or_plan) occur most frequently when students try to form intent. We see similar trends for AI coding during forming input, AI explaining during understand, and AI evaluating during assess step.
Looking more into the ratio of AI used vs. missed, we notice students often missed opportunities to use AI for explanation and evaluation, despite explain code and explain errors being supported by one-click features in Colab notebook. As shown in later analysis (Section 7.1), we suspect students felt overloaded by the information from Gemini.
We also observe that prompt quality could have been largely improved at input step (46% missed opportunities). Instead of clearly instructing the AI (e.g., “write the code to calculate the distribution of the cars by location”) and adding appropriate context (e.g., “help clean up df_games["#total_players"] ”), some students wrote vague prompts such as “fix it” or pasted raw outputs without clear intent in prompt. Without clear_instruction_in_prompt or provide_context_in_prompt, Gemini often produced irrelevant or hallucinated code. This echoes prior work that highlights the importance of teaching students to express their requirements to LLMs (Ma et al., 2025).
Expertise vs. AI usage: Experts use AI more strategically, and have more cognitive budget to explore.
From our logs, we found that expertise got embedded into AI usage in two ways (echoing prior research (Palta et al., 2025)):
  • Technical experts generate code with AI more successfully. As shown in Figure 6A, compared to novices (79%), experts had a 90% success rate in steps where they used AI to generate or edit code. Experts may successfully generate executable code on their first try, or are able to easily debug it. Further analyzing with student grades, it seemed like novice rely more on AI coding being accurate — novice achieving high scores got 82% AI coding success, 4% higher than novices in the low score group (split based on the student homework median grades of 62.2). In contrast, experts achieving high scores actually had slightly lower success on AI coding steps (90% vs. 92%).
  • Technical experts better describe their needs. As shown in Figure 6B, twice as many experts’ prompts were noted to include quality attributes of clear_instruction, decomposed_task, and context. Experts’ prompt was noticeably better, which could result from experts having more domain knowledge (Guo et al., 2024; Palta et al., 2025). Cross-comparing with the high vs. low performing student grade group, we confirmed that prompts noted with high quality attributes were largely contributed by experts who achieved a high score.
Figure 6. (A) Experts achieve a higher step success rate than novices (90% ¿ 79%) when using AI to assist with coding (ai_generate_code and ai_edit_partial_code); Dividing by grades shows more variations. (B) Experts with high grades contributed to most prompts that include high quality attributes.
Figure 7. (A) While experts and novices used AI with similar frequencies, novices were more driven by resolving challenges with AI, and experts were more able to resolve challenges independently without using AI. (B) Compared to novices, experts were more likely to ask AI to explain only when they encountered challenges, and proactively plan and improve even without challenges. These reflect experts’ more strategic distribution of cognitive effort.
We also saw that technical experts using AI more strategically. As shown in Figure 7A, experts and novices used AI at similar frequencies (in around 36% of steps). However, experts were less pressed into using AI when facing challenges — experts’ AI usage (44%) was 10% less than novices for resolving challenges. As shown in Figure 7B, experts ask AI for explanation more often than novices when they ran into challenges (75% vs. 57%); meanwhile, their requests for AI to improve or plan were less tied to challenges than novices (43% ¡ 64%). Experts could be more aware of their goals, or they had more cognitive load to try themselves due to familiarity with the task. Indeed, experts were more capable of resolving challenges without using AI than novices (Figure 7A).
We believe all of these expertise gaps might be closed through training on AI use skills, e.g., when to query AI and with what prompt. For example, a sequence of ai_breakdown_intent $\rightarrow$ ai_generate_code might help a novice obtain more easily digestible code in the first place; if followed by “further simplify your explanation like I’m five”, their chance of correctly generating and understanding AI code can be improved. Similarly, students can start by asking for ai_suggest_steps_or_plans to familiarize themselves with the task, or ai_explain_concepts to understand domain-specific keywords. However, our analysis shows that much of the student grade and effective AI usage still comes down to tech expertise. In the next section, we explore what we should we do better in teaching AI use skills.
7. RQ3: What and how can we train students on AI usage?
The findings in RQ2 suggest that technical experts also use AI more strategically and effectively than novices. Here, we ask: How can we train students to use AI more effectively? What are some skills and learning opportunities for training better AI users?
7.1. The effect of in-class demonstration and extended time-on-task
While we did not experiment with intervention mechanisms for training AI use skills (since the core learning objective was on computational data analysis), we did collect some preliminary signals on what AI use skills might be more easily acquirable than others via just in-class demonstrations and extended time-on-task.
Figure 8. Appropriate AI use ratio pre- to post-instructions. Ratio = $used/(used+missed)$.
Method.
Between HW0 and HW1, the instructor provided a 90-minute live-coding demonstration of the HW0 data cleaning task. The demonstration showcased examples of AI use behaviors in the codebook (Table 3), with explanations of when and why each behavior could be useful. For example, the instructor explained “If you find unfamiliar concepts, you can actually ask Gemini to get you started. We can ask, What are some key steps in Data cleaning?” while typing and sending the actual question and walking through the Gemini’s answer. The demo also covered all the task environment space (e.g., overview of Colab as a product) as well as the AI context space (e.g., the code cell generation has access to the entire notebook, while the chatbot sidebar does not have access to code cell outputs). Students could also re-watch the recording asynchronously. Despite the limitations of lecturing (e.g., less cognitively engaging than active learning and may create an illusion of fluency without fostering durable understanding (Chi and Wylie, 2014; Deslauriers et al., 2019)), this is a lightweight instruction baseline.
In addition to instructions on AI use (and regular lectures on programming and data science domain knowledge), students moved from HW0’s 15-minute time-limited tasks to HW1 and HW2’s bi-weekly assignments (Table 1). Therefore, when we analyze students’ behavioral changes pre- to post-demo, any improvement is a result of both extended time on task and the instruction, and students’ room for improvement builds on top of gains already afforded by extended practice time and baseline instructional exposure.
Figure 9. The distribution of the last successful step in an episode, before and after class demonstration.
Result.
As shown in Figure 9, students were able to sustain progress further along the problem-solving process at post-demo. In HW0, many episodes stalled at early stages (examples of student reported challenges in Table 7; Appendix A) or ended without any successful step (already failed at intent, denoted by “none”). After provided with baseline instructions and ample time, more episodes advanced to later steps of understanding and assessment.
We also computed an appropriate AI use ratio ($used/(used+missed)$) to measure whether students capitalized on available AI opportunities. As shown in Figure 8, appropriate AI use ratios increased post-demo in many behaviors. While not fully mastered or statistically compared, there is a notable progress in students’ prompt quality behaviors by an average of 30% in AI use ratio. Students also showed modest gains in asking ai_explain_concepts. However, evaluation behaviors remained a bottleneck (e.g., post-demo ratios of using AI to critique, asking for code or bug explanations are all ¡ 50%). This suggests a shared hesitation to request or process complex evaluative information and use AI to deepen understanding. These observations hint towards essential but under-taught AI use skills, such as the disposition to persist through trial-and-error, read and interpret long outputs, and refine AI’s outputs (more in Section 7.2).
Note that some AI use behaviors like provide_context_in_prompt are much more frequent (18.06%) than others (Figure 4A), like ai_improve_prompt (0.17%), decompose_task_in_prompt (0.31%), ai_edit_partial_code (0.57%), and ai_breakdown_intent (0.65%). The appropriate AI use ratios of infrequent categories can be less interpretable and easily exaggerated by one or two log instances.
Table 4. The codebook of identified AI use skills by knowledge dimension.
Knowledge Dim. AI Use Learning Objective Description
[Meta]cognitive Monitor when to delegate things to AI and decide what to delegate.
[Conc]eptual Knowing what can be delegated to AI (task space, context space) with calibrated expectations of AI capacity (not under- or over-estimate).
[Conc-Task] Knowing AI Task Space: plan, critique, improve prompt, decompose, explain, edit partial code, etc.
[Conc-Context] Knowing AI Context Space: base Colab Gemini shows no access to closed chats, unreferenced code, imported dataset; different UI (generateCode vs. converse) has different access to context, by the time we ran the study.
[Proc]edural Knowing how to best delegate things to AI for instruction following (clear instruction, context, broken down if overly complex).
[Proc-Clear] Give clear instructions with action verb to AI.
[Proc-Context] Provide correct context to AI without vague pronouns.
[Proc-Decompose] Provide an achievable, decomposed goal for AI instead of asking AI to do too many things at the same time.
[Disp]ositional Willing to actively use AI with trial and error, read and interpret long outputs, and persist through challenging tasks with high intrinsic cognitive load.
7.2. What should we teach on AI use skills
As shown in Figure 8, we observe much room for improvement in AI use skills.To produce actionable insights for instructors, we identified lingering AI use challenges that are not solved by simple classroom demonstrations (using both survey responses and log examples), and summarized concrete learning objectives.
Method
From five post-surveys’ answers to the questions “What were the challenges for you during the tasks when you used GenAI” and “What do you think could help you further improve your performance with GenAI on these tasks?”, we derived learning objectives for AI use skills (Table 4). We adopt the knowledge dimensions taxonomy to categorize AI use skills into conceptual, procedural, dispositional, and metacognitive knowledge (Krathwohl, 2002; Wiggins and McTighe, 2005). Using a similar thematic analysis procedure as described in Section 6.1.2, we iteratively constructed the codebook guided by an open coding process (Strauss and Corbin, 1998), and we validated the codes with our log analysis (Table 6).
Result
In short, effective AI use is not simply reduced to “good prompts” — AI challenges that students encountered often stem from a lack of various AI use skills. As shown in Table 4, we revealed necessary conceptual knowledge (what to know about AI), procedural skills (how to use AI), metacognitive skills (when to apply AI), and dispositional skills (keep engaging with AI). [^5] There can be affective and motivational learning goals, such as self-efficacy and interest to explore. However, these goals are not manifested in our behavioral log data, and we discuss more about possible extensions in future work (Section 8.2).
Table 5. Comparison of expert vs. novice behaviors in resolving KeyError: Column not found. Experts identify missing context (e.g., actual column names) and fix issues using grounded prompts or direct code edits. Novices often rely on vague prompts or trial-and-error, leading to unresolved errors.
Edit on Expert Behavior Novice Behavior
Code KeyError: total_players
 Thinks aloud: ”What’s the actual column name?”
 Checks df.head() output
 Finds and copies #total_players
 Updates code → error resolved
KeyError: price
 Uses “Explain Error” → LLM suggests fixing to Price
 Applies suggestion → resolves error
 Repeats for other columns; some fixes succeed, others (e.g., Interior Color) fail
 Continues blindly copy-pasting → unresolved errors accumulate
 Eventually abandons data cleaning
Prompt KeyError: Current Players
 Prompts: “Column is current_players
 Error persists
 Realizes true name: #current_players
 Clarifies: “Use #current_players to match dataset”
 Model regenerates → error resolved
KeyError: gain
 Prompts: “Don’t use columns that don’t exist” + pastes error
 Thinks aloud: “I already said that, still doesn’t work…”
 Tries variations of prompt → still fails and switches to different EDA question
KeyError: peak_player
 Prompts again: “Don’t use non-existent columns” + pastes error
 Thinks aloud: “I know it doesn’t exist… I don’t know why”
 Gives up and moves on
Differentiate Knowledge Dimensions
Below, we provide some behavioral examples to help illustrate each AI use dimension. However, from our behavioral log data, we could not always identify which knowledge dimension exactly applies. For example, a student may know a particular way of using AI already; however, they may not be able to retrieve it properly at the time they need to use it (available but not accessible knowledge (Tulving and Pearlstone, 1966)). Especially when a student did not use AI, we may not conclude if it was because they did not notice the appropriate time to apply this knowledge (metacognitive), did not have this knowledge stored (conceptual), failed to execute properly (procedural), or lacked the persistence to keep trying (dispositional). Future work may design more targeted assessments corresponding to each type of knowledge in order to differentiate them (Section 8.2).
Metacognitive.
Students often struggled with deciding when to involve the AI and what to delegate. As we analyzed HW1&2 logs (after class instructions), we found multiple instances of students getting stuck in a loop of explain_error and generate_code without re-evaluating their approach (such as the novice’s code edit sequence Table 5). Specifically, these failures stem from repeated KeyError exceptions caused by the AI generating code with nonexistent columns — as Gemini in Colab was not able to retrieve the correct column names from the dataframe, the new code without augmented context kept hallucinating nonexistent columns.
This pattern could illustrate multiple missing AI use skills, such as a lack of metacognitive monitoring: the ability to recognize when a strategy is unproductive and to adjust. The students could avoid delegating AI to do the task that it has demonstrated unreliability and providing the correct column name themselves (such as the expert’s code-edit sequence Table 5), or alternatively ask AI to breakdown the debugging procedure and walk them through a step-by-step process to figure out what is wrong, or add the exact column name as context_in_prompt (such as the expert’s prompt-edit sequence Table 5). Effective metacognitive use would involve pausing to decide, for example, whether to ask the AI for a structured plan (e.g., “list all columns in my dataset”) before persisting with trial-and-error.
Conceptual.
Many failures could stem from the lack of conceptual knowledge regarding the AI’s task space (what it can do) and context space (what information it can access). For task space, most showed improvement in appropriate AI use (Figure 8); however, students may not learn everything they can ask AI to do with simple demonstrations. For instance, the instructor covered using AI to rewrite and improve_prompt, which appropriate use increased from 0% to 100% in post-demo (despite low frequency; Figure 4). Nevertheless, some students still did not realize this use until the later half of the semester: “now I found a smooth way to utilize GenAI – ask it to also re-write my prompt before I use the prompt to generate the code.” (HW3 post-survey). If the demonstration did not covered some AI use, students may develop a wrong assumption of the AI task space. For example, the instructor did not cover ai_edit_partial_code during the demo, and a student demonstrated struggles caused by the lack of this AI use knowledge in their post-survey: “you can’t fine-tune parts of the result; instead, you have to regenerate the whole thing, and sometimes, parts that were correct before end up wrong.”
For context space, the instructor explicitly emphasized that Colab Gemini has limited context space during the demo, with a clear example of generation failure without providing explicit context to the model. However, in students’ logs and post-surveys, we still observe that many students assumed the model could “see” their dataframe and column names without explicit context: “I had to give small details of the code as well to the GenAI, I thought it already has taken the input that is there on the screen.” (post-class survey). As shown in the novice’s prompt-edit sequence in Table 5), a novice might be able to tell that AI is hallucinating non-existent columns, but is still not able to realize they need to augment the context for AI. If the students recall about AI’s limiting context space, they might recognize that the column name needs to be provided in the prompt to avoid hallucination.
Procedural.
Procedural skills concern how to best interact with the AI. Just knowing what AI needs is not enough, as one student noted, “Most of the time, it was difficult to feed context because there were multiple sources to refer to” (post-class survey), reflecting the challenge of providing the appropriate context in a prompt (procedural) even with the conceptual knowledge of AI context space. Similarly, one may know they should break down a complex goal without being able to accomplish it, and end up settling for downgraded goals: “I didn’t really know how to write the prompt for the type of complex data analyses to generate the charts that I wanted, so I had to keep to simple analyses and charts.” In the demo, the instructor exemplified good prompting strategies, such as writing clear_instruction, providing_context, and decomposing_task_in_prompt. However, the post-class behaviors still leave 20% of missing AI use opportunities as room for improvement (Figure 8).
Table 6. Success and failure AI use behaviors with corresponding missing AI use skills.
Success AI use behaviors Failure AI use behaviors AI use challenges due to missing AI use skills
Task Challenge: Form Correct Intent for Task
ai_explain_concepts, ai_suggest_steps_or_plan, ai_breakdown_intent. Stuck as they “don’t know what can I ask AI.”
Did not use AI to assist in forming intent.
[Conc-Task]: know AI can explain concepts, suggest/plan steps, and decompose tasks.
[Proc]: when prompting AI.
[Meta]: monitor when to use AI.
[Disp]: actively explore and persist through ambiguity
Task Challenge: Express Intent in Input (Code)
clear_instruction,
provide_context,
decompose_task_in_prompt.
ai_improve_prompt, ai_generate_code, ai_edit_partial_code.
Fail to clearly specify intent.
Fail to include necessary context $\rightarrow$ hallucinations.
Fail to decompose overly complex intent.
Did not use AI to help formulate the prompt.
Fail to instruct AI to edit code.
[Proc-Clear/Context/Decompose]: use a prompt that clearly specifies intent, contains context, and has a decomposed goal.
[Conc-Context]: know AI needs context augmentation.
[Conc-Task]: know AI can improve prompts and edit partial code.
[Meta]: monitor when to use AI.
[Disp]: patience to re-prompt with more context, instead of repeating the sub-optimal prompt.
Task Challenge: Understand Output (Execution / Text)
ai_explain_code, ai_explain_bug,
Read and interpret AI outputs.
Did not use AI to help explain code.
Did not try to understand the text and just passively copy-pasted.
[Conc-Task]: know AI can explain APIs, code, and bugs.
[Proc]: when prompting AI.
[Meta]: when to use AI.
[Disp]: sustained effort to interpret long outputs.
Task Challenge: Assess Output
ai_critique_output. Did not use AI to evaluate output.
Accepted AI critique without verification.
[Conc-Task]: know AI can critique output.
[Proc]: when prompting AI.
[Meta]: when to use AI.
[Disp]: patience to carefully use AI to evaluate.
Dispositional.
Finally, students may lack persistence and willingness to iterate with the AI and quit from frustration. Some learners disengaged quickly when outputs were erroneous or lengthy. In the afore-mentioned example of repeated KeyError, while AI failed to debug on its own and made repeated errors, some students’ logged actions of generate-code and explain-bug were done within short time intervals of seconds, which indicates that the students did not persist in actually reading or understanding the error explanations or attempting to debug.
The lack of persistence may also result in early stops. For instance, when a student tried to add labels on top of a pie chart, the generated code produced errors, and the fixed code, when simply pressing the “explain error” button, did not fix the issue. The student then stopped trying, accepting the basic pie chart without the label and commented “seems like it either do the label or a pie” (HW2 think-aloud). In comparison, a success case will display resilience, where the student keeps trying to refine outputs until they achieve the intended goals. For example, as shown in a sequence of prompt iterations, the student persisted through different AI mistakes to produce the ideal visualization for EDA:
Given the above dataset, I need to see 5 years of pie charts comparing genre to release date. The goal is to understand which genres, as a percentage of all genres, were released each year
$\rightarrow$ [copied: … df_games.info() ] Regenerate the ”genre to year” comparison again … All operations should be done within this dataset for all prompts going forward
$\rightarrow$ The graphic is too noisy to glean anything useful. Can we pick the top n games … if a game has less than 3% share … we don’t display it …?
$\rightarrow$ You did the opposite of what I asked. … display the genres OVER a certain threshold. 3% was an example. Can you suggest a better way to find this statistically important threshold and then re output using it?
$\rightarrow$ Pie chart is not the answer. Can we try a histogram or bar chart? If those aren’t wise, what would you recommend for comparing the genre to total players by year?
$\rightarrow$ regenerate to make the grouped bar chart
$\rightarrow$ Change this to output one such chart per year. There should be one chart for every year in the range
(HW2 log).
8. Discussion and Future Work
8.1. Design and Pedagogy Implications
Make AI mental models explicit.
Unlike simple chat interfaces, AI integrated in a notebook exposes multiple affordances (code cells, outputs, error panes, menu buttons, etc.). We observed in the study that students transferred assumptions from more capable general-purpose AIs to their Colab working environment, miscalibrating what the embedded model could access or do. This produced confusion about AI task space (what to ask) and context space (what the model can “see”). As shown in Table 5, a novice during think-aloud prompted the model “Don’t use nonexistent columns, use what you have in the datasheet” after three repeated KeyError and commented “Perplexity can do that,” assuming that Colab Gemini would have access to the imported dataframe like other systems. It took them more than five repeated failures to realize that they should adjust their mental model for the system at hand, highlighting the need for explicit conceptual models and metacognitive updating abilities. On the design side, systems with GenAI assistants should make their task and context scope more legible (e.g., what files and variables are visible to GenAI, what operations are supported, etc). Teaching “what to ask this specific AI” and “what this specific AI knows” is likely getting obsolete as models and systems rapidly evolve. More importantly, instructions are needed for metacognitive abilities to quickly adjust mental models to different AI system capabilities.
Facilitate evaluation by providing less information and scaffolding verification.
As we found in Section 7.1, using AI to gain additional, evaluative information was a consistent challenge. Asking AI to explain code, bugs, or critique outputs remained limited, despite its potential benefit to improve understanding. We especially observe that students may not persist in reading convoluted output, reflecting reluctance or difficulty in engaging with evaluative or complex information. While there are one-click affordances (explain error buttons) that aim to make it easier to start evaluation, the cognitive load of seeing a lengthy error explanation often creates resistance or reduces the chance for students to learn critical evaluation skills.
On one hand, we need instructions on dispositional skills to help humans stick to evaluation and mitigate the risk of overreliance. On the other hand, in terms of feedback design for novice students or users, less is more. Explicit instructions that focus exclusively on evaluating LLM outputs may alleviate this cognitive load in a complex task environment (Ma et al., 2024). Additionally, it is possible to lean on the rich log data to trigger real-time hints and output style changes. For example, repeated “fix it” after KeyError in short intervals suggests struggles of evaluation, so instead of providing a full code fix saying “If your column name is X, the code will look like Y [lengthy code] [lengthy explanation],” generating a diagnosis plan or hint for the user one step at a time might be more effective, e.g., “Run df.columns and check the output. Copy-paste the exact column name you want to use in code.”
In terms of design, generating UIs (Chen et al., 2025a) instead of lengthy text may help alleviate the evaluation burden. Designs that intentionally insert intermediate checkpoints can help decompose big evaluation plans into more manageable chunks, or frontload evaluation by asking users to provide requirements before generation, e.g., in the form of plan skeletons and verification checklists (Kazemitabaar et al., 2024a). Pedagogy can mirror this by requiring learners to specify success criteria and requirements before requesting code (Ma et al., 2025).
8.2. Limitations and Future Work
Going deeper: ground in assessments and controlled experiment to study factors influencing effective LLM use
Our analyses suggest that students’ technical expertise significantly predicts performance during LLM use (similar to prior literature (Chen et al., 2024)), whereas LLM and communication expertise factors did not (Section 5). Other work also shows that LLMs can narrow gaps between novices and experts (Pickering et al., 2025), or even impede expert performance under certain conditions (Becker et al., 2025). This highlights that technical expertise is not the only factor shaping effective LLM use. Our log analyses revealed the importance of AI use strategies, which point toward the need to define and evaluate “AI use expertise” more directly, as what we have attempted in Section 7.2.
To probe the interplay between technical expertise and LLM use, we compared performance between experts and novices in the timed HW0 under conditions with and without LLMs (Table 1). A Mann–Whitney U test (as data violated normality assumptions) (McKnight and Najab, 2010) showed a significant expertise gap when students did not use LLMs ($p=0.027<0.05$), but the gap disappeared when LLM was used ($p>0.05$). This suggests that under time pressure, LLMs may temporarily offset expertise advantages, indicating that factors such as time constraints also influence the effectiveness of LLM use. We also measured affective factors (motivation, self-efficacy) in self-reported survey items (Appendix B), but none significantly predicted performance.
Our sample included 36 students in authentic classroom contexts. While this ensured ecological validity, it is a small sample and constrained our ability to isolate factors such as undisclosed external tool use (e.g., ChatGPT alongside Colab Gemini) or overlapping expertise dimensions. For example, Python and data science proficiency were moderately correlated in our sample, so we could not examine their independent effects on effective LLM use. Future studies should recruit distinct participant groups or employ pretests. More robust performance-based assessment and controlled experiments are needed to disentangle the interplay between more nuanced factors such as domain expertise, implementation expertise, AI-use expertise, affect, and situational factors like time pressure.
Going broader: extend to more diverse task contexts.
The tasks employed in our study were appropriate for classroom learning but were not “LLM-hard:” current models could achieve high accuracy. Future work could incorporate more complex tasks that better reveal the value of AI use skills and enable observations on how learners adapt when models struggle.
While the learning objectives for AI use skills (Table 4) were derived from programmatic data analysis tasks, they should extend beyond our study context. Many of the objectives are broadly relevant to problem-solving and higher-order thinking skills; for instance, clarifying the task space of any tool, communicating requirements effectively with decomposition, and verifying results. Future research could examine whether instruction around AI use transfers to these more general competencies, and explore how to evaluate AI use expertise more systematically (as self-rated LLM expertise may not be appropriate).
Finally, we chose Google Colab for the class because it embeds a free, in-house Gemini with different UIs, which enabled us to explore diverse behavioral AI use patterns in our research questions. However, our data collection relied on a custom logging script in the browser, which can be less robust than existing Jupyter notebook logging toolkits like JELAI (Valle Torre et al., 2025). For example, our approach may capture an auto-completion as users’ code edits (although infrequent), and Colab’s frequently changing interface raises concerns about long-term replicability. Future work could build on established logging toolkits to improve data quality and generalizability.
9. Conclusion
LLMs promise to broaden access to technical work, yet our study shows that not everyone benefits equally. Technical expertise remained the strongest predictor of success in homework, and experts engaged with AI more strategically than novices. Behavioral log analysis highlighted that experts provide clearer context, decompose tasks, and strategically use AI, while novices tend to rely on AI reactively in the face of challenges. While demonstrations and extended time helped students adopt clearer prompting practices, evaluative skills remained challenging. These findings suggest that effective AI use is not just about producing good prompts, but also more about cultivating durable competencies in multiple dimensions. We translate these insights into design and pedagogy implications for AI literacy: explicate mental models, clarify before generating, and scaffold verification. Ultimately, effective AI use is not just better prompts — it is about better problem solving.
Acknowledgements.
Thanks to all the participants for the study, and thanks to all LearnLab members who provided feedback on this work. Thanks to the National Science Foundation (award CNS-2213791, 2414915) and Google Academic Research Award for partial support of this work. Acknowledgment of LLM Use. We acknowledge the use of ChatGPT to assist with language editing, summarization, and restructuring of text during manuscript preparation. Claude was used to assist the analysis of log data; implementation details are included in Section 6.1.2, prompts and settings are attached in Appendix D. Copilot was used to assist the visualization and code writing to process and clean our dataset, and Gemini in Google Colab was used to prepare the homework materials and reference solutions.
Appendix A Codebook for task challenge and AI use success and failure behaviors
Norman’s Stages of Action (Tenner, 2015) describes the user cognitive process as a sequence of forming the goal and the intention, specifying and executing an action, perceiving and interpreting the world state, and evaluating the outcome. Inspired by this theoretical framework, we thematically coded the survey answers for “What were the challenges for you during the tasks when you did not use GenAI?” from the HW0 post-survey, focusing on task challenges that exist even without AI use and could be addressed with effective AI use.
As shown in Table 7, students reported challenges throughout forming intent, writing code, understanding output, and assessing output. For example, students reported that they “did not know where to start at all” (form intent), had challenges in “putting the right code and iteratively coding on it” (write code), “interpreting the errors” (understand output), and “knowing what kind of answer is good” (assess output). Students who did not have any programming or data science background reported that they “do not know what to do” and “have no idea if I’m doing it right or even understand what I’m doing.” Even students with some programming experiences described the process as “trying to build a house with just a hammer and some nails” and “working blindfolded in a room full of puzzle pieces”.
Table 7. Processes and task challenges in workflows. Effective task completion is a combination of both correctness and efficiency.
Process Process Description Task Challenge Task Challenge Description
Form
Input
Expressing intent through correct Python code & producing desirable output for data science tasks Form Correct Intent for Task Effectively knowing what to do (e.g., edit code, refine visual, clarify concept, break down complex goal, etc.).
Express Intent in Code Effectively generate code (for the data science task) for the right intent to produce desirable outputs.
Process Output Making sense of code output; checking correctness and alignment with goals Understand Output (Execution / Text) Effectively read and understand code execution output, such as error messages, displayed images, tables, and text.
Assess Output Effectively evaluating whether the output is correct and aligned with task goals.
Appendix B Survey Instruments
Here, we summarize survey instruments administered during the study. Questions are grouped into (i) shared items common across multiple surveys, and (ii) unique items included only in the Pre Survey, HW0, HW1–HW4, or Final Post Survey. Some survey items are drawn from existing measures such as NASA Task Load Index (TLX) for cognitive load (Hart and Staveland, 1988).
B.1. Pre Survey Only Questions
  • Demographics: “How old are you?”; “What is your gender?” (Male/Female/Non-binary/Prefer not to say).
  • Background: “What major did you study for undergraduate/college?”; “What program are you studying now at CMU?”.
  • Expertise: “How would you assess your level of expertise in the following areas?” (Programming, Communication, GenAI/LLM, Data cleaning, EDA, ML, Storytelling, Car industry [dataset domain], Video game [dataset domain]; 5-point scale).
  • Interest: “How would you assess your level of interest in the following areas?” (same domains, 5-point scale).
  • Frequency: “How frequently do you use GenAI/LLM for Programming, Communication, Data cleaning, EDA, ML, Storytelling?” (Never–Daily).
  • Experience: “How many years of experience do you have with Excel, R, Python, Jupyter/Colab, pandas, Altair, scikit-learn?” (0, 1–2, 3–5, 6–10, ¿10).
  • Confidence: “How confident are you in your ability to do the following tasks on your own?” (Programming, Data cleaning, EDA, ML, Storytelling).
  • GenAI ability: “How confident are you in GenAI/LLM’s ability to do the following tasks?” (same domains, 5-point scale).
  • Attitudes: “I am confident in my ability to communicate to GenAI/LLM so that it does what I want.” (5-point Likert).
  • Expectations: “What do you want to learn from this class?”.
B.2. Shared Questions Across Post Surveys
  • GenAI tool usage. “What GenAI did you use for this homework? (Colab Gemini, ChatGPT, Copilot, Perplexity, Claude)”.
  • Perceived difficulty. “How mentally demanding/difficult was the task?” (5-point Likert from “Not difficult at all” to “Extremely difficult”).
  • Effort. “How hard did you have to work for the task?” (5-point Likert from “I do not need to work hard at all” to “I need to work extremely hard”).
  • Frustration. “How irritated, stressed, and annoyed did you feel during the task?” (5-point Likert).
  • Interest. “How interesting/engaging was the task?” (5-point Likert).
  • Self-evaluation. “How successful were you in performing the task? How satisfied were you with your performance?” (5-point Likert from “Not successful/satisfied at all” to “Extremely successful/satisfied”).
  • Confidence. “How confident are you that you did well on the task?” (5-point Likert).
  • Peer help. “How much did you ask your peer/classmate for help with the task?” (5-point Likert from “Not at all” to “All the time”).
  • Open-ended reflections.
    • “What was challenging for you during the tasks when you used GenAI?”.
    • “How did you use GenAI during the tasks?”.
    • “Did you find using GenAI helpful during the tasks? Why or why not?”.
    • “What do you think could help you further improve your performance with GenAI on these tasks?”.
    • “What help did you offer to or receive from peers/classmates? What questions do you ask them instead of Google or GenAI?”.
B.3. HW0 Post Survey Only Questions
  • Task assignment: “Which tasks did you do with GenAI?”; “Which tasks did you do without GenAI?”.
  • Link: “Link to Zoom recording”.
B.4. HW1–HW4 Post Survey Only Questions
  • Each survey contextualized task focus:
    • HW1: “Part 1. Data cleaning”.
    • HW2: “Part 2. Exploratory data analysis”.
    • HW3: “Part 3. Machine learning”.
    • HW4: “Part 4. Data-driven storytelling”.
B.5. Final Post Survey Only Questions
  • Retrospective self-assessment of expertise, interest, and confidence in self vs. GenAI ability across domains: E.g., “How would you assess your level of expertise in the following areas?” (same as Pre Survey, 5-point scale).
  • Trust, communication, and AI self-efficacy statements: E.g., ‘Generally, I trust GenAI/LLM” (same as Pre Survey).
  • Open-ended retrospective variants:
    • “What was challenging for you during the tasks when you used GenAI? Any similarities or differences across the assignments?”.
    • “How did you use GenAI during the tasks? Any similarities or differences?”.
    • “What did you learn from this class? What do you wish that you could learn more about?”.
Appendix C Composite Competency Inclusions
Technical competency (composite of programming, data science, and tools)
  • How would you assess your level of expertise in the following areas: [Programming; Data cleaning; Exploratory data analysis; Machine learning; Data-driven storytelling]
    Scale: Novice (1); Beginner (2); Intermediate (3); Advanced (4); Expert (5)
  • How many years of experience do you have in using the following tools? [Python; Excel; Jupyter/Colab notebook; pandas; altair; scikit-learn]
    Scale: 0 (1); 1–2 years (2); 3–5 years (3); 6–10 years (4); ¿10 years (5)
  • How confident are you in your ability to do the following tasks on your own? [Programming; Data cleaning; Exploratory data analysis; Machine learning; Data-driven storytelling]
    Scale: Not confident at all (1) – Extremely confident (5)
LLM/LLM-communication competency (combined)
  • How would you assess your level of expertise in the following areas: [GenAI/LLM]
    Scale: Novice (1); Beginner (2); Intermediate (3); Advanced (4); Expert (5)
  • How frequently do you use GenAI/LLM for: [In general]
    Scale: Never (1); Rarely (2); Monthly (3); Weekly (4); Daily (5)
  • Please rate the following in terms of how much you agree or disagree with each statement: [I am confident in my ability to use GenAI/LLMs; I am able to have LLMs complete the tasks I instruct; I am confident in my ability to communicate to LLM so that it does what I want; Generally, I trust GenAI/LLM]
    Scale: Strongly disagree (1) – Strongly agree (5)
Communication competency
  • How would you assess your level of expertise in the following areas: [Communication]
    Scale: Novice (1); Beginner (2); Intermediate (3); Advanced (4); Expert (5)
  • Please rate the following in terms of how much you agree or disagree with each statement: [I am confident in my communication ability in general]
    Scale: Not confident at all (1) – Extremely confident (5)
Appendix D Prompt for Log Analysis
A temperature of 0 is used with the model claude-sonnet-4-20250514 and max_tokens of 20000, in the following two steps as a Chain-of-Thought:
D.1. Segmentation Prompt
Task
Turn a chronological JSON log of events from Google Colab (students completing a data-science task with help from the Gemini model) into episode segmentation.
Key Definitions:
Episode
: A contiguous sequence of actions (form intent, form input, understand output, assess output) driven by a single overall intent (e.g., “remove duplicates in data column X”, “fit a regression model to predict Y,” “creating visualization to analyze the trend of Z”). An episode ends when the student succeeds, gives up, or pivots intent.
Decision boundary for different intent (and split into new episode): a different EDA question is a different intent a different plot is a different intent, even if it visualizes the same relationship or answers the same EDA question cleaning a different column is a different intent removing outliers is different intent from making a column dtype numeric, even if they are both cleaning the same column Episode Steps: each episode includes the following four steps; early stops/failures may not reach beyond step 1: (1) Form Intent: Deciding what needs to be done to progress the task (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count, clarify concept, break down a complex goal). Each unique intent should start a new episode. Writing comments and prompting model are common behaviors that can be used to infer intent. (2) Form Input: Translating intent into input (Python code in edit_code or an AI prompt in converse or generateCode) that can generate the desired result. (3) Understand Output: Reading and interpreting the system’s response (llm output in event_response, executed_output, error messages in executed_error, plots, tables, or text). (4) Assess Output: Evaluating whether the output is correct, useful, and aligned with the intended goal. Instructions:
1. Segment logs into episodes:
Start a new episode when you observe a clear change in intent (see decision boundary above), OR if there is a time gap $\geq$ 15 minutes between consecutive events (use datetime_timestamp). Ensure all log events are covered. Episodes should be consecutive, from the first log to the last. Prefer fewer coherent episodes unless there is a clear goal shift or long gap. Each episode must have timestamp_start and timestamp_end taken from the episode’s first/last event.
2. Identify the episode’s overall intent: Infer from code, prompts, and outputs (e.g., remove outliers in a column, create a bar chart, refine a visual by sorting by count). Write 1–2 sentences describing what the user was trying to do and what happened, inferred from the log events. Output Format (JSON array of episode objects):
  "episode_id": 1,
  "timestamp_start": "datetime_timestamp",
  "timestamp_end": "datetime_timestamp",
  "events": ["evt_0", "evt_1", ... all event_ids in the episode],
  "overall_intent": "inferred user intent from logs"
D.2. Step-wise Annotation Prompt
Task
Turn an episode of chronological JSON log events from Google Colab (students completing a data-science task with help from the Gemini model) into episode steps with fine-grained actions and returns. Then annotate AI usage and its success/failure in the episode.
Input
Each event contains fields like: event_id, event_name, event_category, event_content, event_response, and datetime_timestamp. AI use behaviors are recorded in events with event_type = converse or generateCode.
Key Definitions
There are four different types of steps in an episode:
(1) Intent – forming a correct goal (2) Input – expressing the goal clearly in code or prompt to generate desired output (3) Understand – interpreting output/errors (4) Assess – judging output correctness & alignment with goal
Users may display different AI-use behaviors in different steps of the episode: intent-step: asks AI to explain concepts or suggest steps; asks AI to break down steps. ai_explain_concepts ai_suggest_steps_or_plan ai_breakdown_intent input-step: asks AI to improve/clarify the prompt; ask AI to edit code or generate code ai_improve_prompt ai_edit_partial_code ai_generate_code understand-step: asks AI to explain code or error, press_explain_error, press_explain_code, or clarifies APIs ai_explain_bug_or_error ai_explain_code_or_api assess-step: asks AI to critique output; queries alignment with goals/insight quality ai_critique_output step-agnostic-ai-use: good AI use can happen across steps (whenever the user is using AI) clear_instruction_in_prompt provide_context_in_prompt decompose_task_in_prompt Behavior Examples
The user may encounter challenges at any step in the episode.
Some example behaviors that demonstrated challenges in each step: intent-step: user stuck (“don’t know what to do”), failure to form intent (e.g., “what else can I do?” / “give me steps”), wrong intent (e.g., asks for adding “\$” instead of making a price column numeric), cannot decompose a complex goal. input-step: vague pronouns (“fix it”) without shared context, missing/incorrect context leading to hallucinations (e.g., KeyError of made-up columns), faulty code (syntax error, wrong API usage), edit_code followed by executed_error. understand-step: copy-paste generated code blindly, immediate “fix it” after press_explain_error without reading, repeating error loops with no adjustment. assess-step: accepts generation without verification (very short intervals), keeps ineffective chart/insight misaligned with goals. Some success behaviors for each step: intent-step: final intent is clear and correct (e.g., valid cleaning procedure, valid EDA question) input-step: input code/prompt are clear and correct (errors resolved) understand-step: demonstrated understanding via a concrete next action or time spent reading assess-step: result (data/chart/stats) is correct; or student requests targeted improvement based on assessment Instructions
1. For each step of the episode, trace key events and summarize into Actions & Returns with evidence:
Action: initiate an attempt (e.g., edit code, enter prompt, send Gemini prompt, UI button like execute after editing code, press_explain_error, etc.) Return: immediate observable system feedback (stdout/stderr, exceptions, plots, metrics, AI response) Steps and evidence event_ids must be in chronological order.
2. For each step, annotate if a challenge was encountered, and whether AI was used strategically to overcome it: Use the challenge/success behavior examples to mark encountered_challenge. If AI was used well, tag behaviors (e.g., ai_explain_concepts, clear_instruction_in_prompt) and briefly explain how it helped. If not overcome, note which AI-use behaviors could have helped and why. 3. Decide if the step was successful, and explain: step_success can be true/false (partial progress or step not reached counts as false). Provide 2–3 sentences citing how the outcome is inferred from logs; include representative event_names and short excerpts from event_content/errors/prompts or relevant timestamp gaps. Steps can be independent (e.g., wrong intent but correct understanding/assessment). Visual-output assessment
If the output involves a chart, evaluate the output code to judge whether the visualization answers the EDA question:
(1) Identify if there is visualization code (e.g., alt.Chart) intended to answer an EDA question. Evaluate the code even if the actual image is not available. (2) Decide if the visualization is appropriate. If not, did the user refine it appropriately? Examples of assessment failures include: accepting an overly-busy chart (e.g., all genres when only top genres matter); using the wrong plot type (e.g., line instead of scatter for correlation); using a boxplot to show average rating (producing short lines rather than showing a range).
Output format
Return output strictly as a JSON array of objects with these fields only:
  "step_id": 1,
  "step": "intent | input | understand | assess",
  "actions": [
    {"action": "event_name + short content excerpt", "summary": "what happened",
     "evidence": "evt_0" | ["evt_0","evt_1"]}
  "returns": [
    {"return": "stdout|stderr|exception|plot|ai_response", "excerpt": "string",
     "evidence": "evt_1" | ["evt_0","evt_1"]}
  "encountered_challenge": true | false,
  "ai_used_to_overcome_challenge": [] | ["ai_explain_concepts","clear_instruction_in_prompt"],
  "how_ai_helped": "" | "briefly explain ai_used_to_overcome_challenge",
  "ai_usage_that_couldve_helped": [] | ["provide_context_in_prompt"],
  "why_ai_could_help": "" | "briefly explain ai_usage_that_couldve_helped",
  "step_success": true | false,
  "reasoning": "how is the step & outcome inferred from logs"